what is the answers to this eqaution Im not sure if i have the right answer
3 −  5/2 x − 3x²   =   0

Answer by AnlytcPhil(1806)

what is the answers to this eqaution Im not sure if i have the right answer
3 − 5/2 x − 3x² = 0
Clear of fractions by multiplying through by 2
6 - 5x - 6x² = 0
Rearrange in descending powers:
-6x² - 5x + 6 = 0
Multiply through by -1 to remove initial negative sign:
6x² + 5x - 6 = 0
Factor left side:
(3x-2)(2x+3) = 0
3x-2 = 0 ; 2x+3 = 0
3x = 2 ; 2x = -3
x = 2/3; x = -3/2
Edwin J